Sve & 10.00 points Problem 17-16 Relation of rights to EPS and the price earnings ratio [LO17-3 The current market price of Waker common stock is $88 per share rights-on. The companys net income this year is $26.00 million. A rights offering has been announced in which 720.000 new shares will be sold at $00 50 per share. The subscription price plus mine rignts is needed to buy ane of the new shares (Do not round intermediate caleulations and round your answers after the rights offering? (Assume thare is no change in and round your O Type here to search
M Chapter 17 HW b. What would the the market value of the stock, except for the change when the stock begins trading ex-rights.) (Do not round after the nghts offering?(Assmethere rochange in intermediate calculations and round your answers to 2 decimal places.) Problem 17-16 Relation of earnings ratio [LO17-3] first option to purchase new shares Type here to search
Answer #1

Answer 1:

Net Income = $26.00 million

Number of common stock outstanding = 7.2 million

Earnings Per share = (Net Income - Preference dividends) / Number of common stock outstanding

= $26 / 7.2

= 3.61111

= $3.61

Price Earning ratio = Market Value per Share / Earnings per Share = $86 / ($26 / 7.2)

= 23.81538

= 23.82

Earning Per share$3.61 Price Earning ratio 23.82

Answer 2:

After rights offering number of common outstanding = 7,200,000 + 720,000 = 7,920,000

Earnings Per share = $26,000,000 / 7,920,000 = 3.28283 = $3.28

Ex rights price of share = (Market Value of shares prior to rights issue + Cash raised from rights issue) / Number of shares after rights issue

= ($86 * 7,200,000 + $80.50 * 720,000) / (7,200,000 + 720,000) = $85.50


Price Earning ratio = $85.50 / ($26,000,000 / 7.920,000​​​​​​​) = 26.044615 = 26.04

Earning Per share$3.28 Price Earning ratio 26.04